# Quantifying the noise sensitivity of the Wasserstein metric for images

### v1

@@ 3.3 Multiscale $W_{p}$ bound on a single image and its noisy version
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions, and let $\varepsilon$ satisfy Assumption 1. Then
$\mathbb{E}W^{\pm}_{1}(\mu,\mu+\varepsilon)=\mathbb{E}W_{1}(\varepsilon^{+},\varepsilon^{-})\leq\frac{2}{\sqrt{\pi}}\sigma n\log_{2}n+\frac{1}{2\sqrt{\pi}}\sigma n.$ (15)
@@
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$. Let $\epsilon$} be a signed measure on the grid that satisfies Assumption 1. For convenience, we again assume that $n$ is a power-of-two. Then, for $p>1$ with $p\in\mathbb{N}$,
$\mathbb{E}\left[\left(W_{p}^{\pm}(\mu+\epsilon,\mu)\right)^{p}\right]\leq\frac{4}{\sqrt{\pi}}n\sigma.$ (16)
@@
$\mathbb{E}\left[W_{p}^{\pm}(\mu+\epsilon,\mu)\right]\leq\left(\frac{4}{\sqrt{\pi}}n\sigma\right)^{1/p}.$
@@ 3.4 Multiscale $W_{p}$ bound on two images and their noisy counterparts
Let $\mu,\nu:G_{n}\to[0,1]$ be two probability measures on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let ${\epsilon}_{\mu},{\epsilon}_{\nu}:G_{n}\to\mathbb{R}$ be signed noise measures that satisfy Assumption 1. For convenience we assume that $n$ is a power-of-two. Then
$\mathbb{E}\left[W^{\pm}_{1}(\mu+{\epsilon}_{\mu},\nu+{\epsilon}_{\nu})-W^{\pm}_{1}(\mu,\nu)\right]\leq\frac{4n\log_{2}n+n}{\sqrt{\pi}}\sigma+\frac{\sqrt{2}}{n}.$ (18)
Even though the Wasserstein 2-distance is often used in applications and has nice theoretical properties in the continuous setting —such as the Brenier–McCann theorem (Brenier, 1991), its signed counterpart does not enjoy the same metric properties as the signed Wasserstein 1-distance, as was already hinted at in the introduction.
@@
Let $\mu,\nu:G_{n}\to[0,1]$ be two probability measures on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let ${\epsilon}_{\mu},{\epsilon}_{\nu}:G_{n}\to\mathbb{R}$ be signed noise measures that satisfy Assumption 1. For convenience we assume that $n$ is a power-of-two. Then
$\mathbb{E}\big[W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})\big]\leq\left(\frac{\sqrt{2}}{2}\right)^{1-\frac{1}{p}}W_{1}(\mu,\nu)^{\frac{1}{p}}+\frac{\sqrt{2}}{2}\bigg(\frac{4}{\sqrt{\pi}}n\log_{2}n+\frac{2}{\sqrt{\pi}}n\bigg)^{\frac{1}{p}}\sigma^{\frac{1}{p}}.$ (19)
@@ 4.1 Quantitative validation of noise scaling
The first experiment we conduct aims to quantitatively measure how the distance between an image and its noisy counterpart scales when increasing noise variance. This allows for a direct comparison between the empirical behavior of each metric and the theoretical scaling laws derived in Theorem 3. The results are reported as Figure 2. All transport costs calculated are exact and were calculated by using the POT python package Flamary et al. (2021)
Figure 2: $W_{1},W_{2}$ and $L^{2}$ (markers) distances plotted against their fits and their theoretical bounds (dashed lines)
To this end, we performed 100 independent trials, each time selecting a new, random 32x32 pixel image from the DOTMark 1.0 MicroscopyImages dataset (Schrieber et al., 2017). For each image $\mu$, we generated a noisy version $\mu+\epsilon$ by adding zero-sum noise $\epsilon$} satisfying Assumption 1, with variances ranging from $10^{-7}$ to $1$. We computed the difference between the original image and the noisy one for $L^{2}$, $W_{2}$ and $W_{1}$ identifying the image with the torus. This empirical result, where the $W_{2}$ distance scales with an exponent of approximately 0.5, suggests that the bound derived in Theorem 3 correctly captures the behavior of the signed 2-Wasserstein as a function of the noise variability $\sigma$.
@@ 4.2 Visualizing robustness of inter-image distances
We now investigate how well the different metrics preserve the original distance between two images when the latter are progressively corrupted by noise. For this experiment, we selected two distinct $32\times 32$ pixel images from the DOTMark dataset and simultaneously corrupted them with different instances of zero-sum additive noise with a standard deviation ranging from $10^{-4}$ to $10^{-2}$. At each noise level, we computed the $W_{1},W_{2}$ and $L^{2}$ distances between the two noisy images. The results were averaged across 100 experiments. To evaluate stability, we computed a “Distance Ratio” by dividing the distance between the noisy images by the constant distance between the original, clean images. A ratio that remains close to 1 indicates that the metric’s measurement reflects the underlying signals rather than the noise. The output is displayed on Figure 1, which we already exhibited in the introduction.
Figure 3: Distances between two randomly sampled images from the DOTMark microscopy dataset, both being corrupted with noise sampled from the zero-sum normal distribution, in dashed (matching colors) we have the bounds for each $p$ from Theorem 3.
@@
We repeated the distance ratio experiment from the previous section on four distinct image classes: white noise, typical cryo-EM projections, classic microscopy images, and synthetic images of two widely separated squares. These classes represent types of images which are ranging from pure noise, to having all the mass centered in a single pixel. In this experiment we aim to show how $W_{2}$ scales favorably throughout the whole range.
4.3 Application to Cryo-EM image alignment
For the final set of experiments, we have taken 20 different projections of the hsp90 protein in different conformational states. The image data was generated with cryojax (CryoJAX Developers, 2025), relying on the structure from the Protein Data Bank, which is based on the work of Shiau et al. (2006). The location of the protein was shifted and the pictures were corrupted with noise with $\sigma=0.01$. Remark that the shift and very poor signal-to-noise ratio are not whimsical but accurately depict the challenges in the field of cryo-EM.
The goal is to assess how well each metric can recover known geometric relationships between particle images that undergo rotation and translation, all under a heavy-noise regime typical of real-world data.
To illustrate the difficulty of the task, Figure 5 shows a sample of the original, clean images alongside their noisy counterparts. The noise level is high enough to make the underlying molecule almost completely unrecognizable by eye, simulating the low signal-to-noise ratio of raw micrographs.
Figure 5: Original images of the molecule on the left, noisy images with $\sigma=10^{-2}$ of the molecule on the right.
@@ 4.3 Application to Cryo-EM image alignment
Figure 6: The left image of each pair corresponds to the distance matrix between the different structures in the absence of noise, while the right represent averaged distances over 100 experiments where the images have been corrupted by random noise. Each pair of pictures corresponds to a different metric. A more gradual blue-to-red gradient better captures the underlying motion, suggesting that $W_{2}$ performs best while being robust.
@@ 4.4 The decreasing distance phenomenon
Figure 7: Distances between two images, one has a mass of 1 in (8,8) and the other in (24,24) and zero everywhere else, both being corrupted with noise sampled from the zero-sum normal distribution, in dashed (matching colors) we have the bounds for each $p$ from Theorem 5.
@@
Figure 8: Top row: Source image alongside images of where the mass of the original pixel (8,8) goes when the target and source images are corrupted by noise. Bottom row: Target image alongside images of where the mass of the target pixel (24,24) comes from when the target and source images are corrupted by noise.
@@ 5 Conclusion and future work
Despite these results, there remains venue for additional work. A primary challenge would be to establish a (sharp) bound for $\mathbb{E}W_{p}^{\pm}(\mu_{\epsilon},\nu_{\epsilon})-\mathbb{E}W_{p}^{\pm}(\mu,\nu)$, which is hindered by the lack of triangle inequality. The numerical experiments further suggest that our bounds, despite capturing the correct behavior, are not tight. Finally, as our theory suggests that robustness increases with higher values of $p$, the interest of such choices of exponents for practical applications should be investigated in future works.
Appendix A Deferred proofs
@@ A.1 Proof of Propositions
Proposition 3 (Wasserstein Distance Decomposition).
Proof of Proposition 3.
$m(A)\leq\mu(A)\text{ and }m(A)\leq\nu(A).$
The remaining, disjoint parts of each measure are respectively $\mu^{\prime}:=\mu-m=(\mu-\nu)_{+}$ as well as $\nu^{\prime}:=\nu-m=(\nu-\mu)_{+}$. Thus, we can write:
$\mu=m+\mu^{\prime}\qquad\nu=m+\nu^{\prime}$ (21)
$W_{p}^{p}(\mu,\nu)\leq W_{p}^{p}(\mu^{\prime},\nu^{\prime})$ (23)
Proposition 4.
Let $\mu,\nu:G_{n}\to[0,\infty)$ be images on the square grid $G_{n}$ with spacing $h=1/n$, and let $\epsilon_{\mu},\epsilon_{\nu}$ satisfy Assumption 1. Identifying $G_{n}$ with the $2$-torus, let $D:=\mathrm{diam}(G_{n})=\sqrt{2}/2$. Then, for any $p\geq 1$,
$W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})\leq D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big)^{\frac{1}{p}},\qquad\epsilon^{*}:=\epsilon_{\mu}-\epsilon_{\nu}$ (24)
Proof.
By definition of the signed distance,
$W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})=W_{p}\big((\mu{+}\epsilon_{\mu})_{+}+(\nu{+}\epsilon_{\nu})_{-},\ (\nu{+}\epsilon_{\nu})_{+}+(\mu{+}\epsilon_{\mu})_{-}\big).$ (25)
$W_{p}^{\pm}(\mu+\epsilon_{\mu},\nu+\epsilon_{\nu})\leq W_{p}\big([\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{+},\ [\nu{+}\epsilon_{\nu}{-}\mu{-}\epsilon_{\mu}]_{+}\big).$ (26)
Applying this to $(\alpha,\beta)=([\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{+},[\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{-})$ yields
$W_{p}([\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{+},[\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{-})\leq D^{1-\frac{1}{p}}\bigg(W_{1}([\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{+},[\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{-})\bigg)^{\frac{1}{p}}.$ (28)
Using Proposition 6, we conclude
$D^{1-\frac{1}{p}}\bigg(W_{1}([\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{+},[\mu{+}\epsilon_{\mu}{-}\nu{-}\epsilon_{\nu}]_{-})\bigg)^{\frac{1}{p}}$
$\qquad\leq D^{1-\frac{1}{p}}\bigg(W_{1}(\mu,\nu)+W_{1}\big((\epsilon_{\mu}-\epsilon_{\nu})_{+},(\epsilon_{\mu}-\epsilon_{\nu})_{-}\big)\bigg)^{\frac{1}{p}}.$
Since both $\epsilon_{\mu}$ and $\epsilon_{\nu}$ are normally distributed, we can say that $\epsilon^{*}:=\epsilon_{\mu}-\epsilon_{\nu}$ is also normally distributed, with $\operatorname{cov}(\epsilon^{*})=2\operatorname{cov}(\epsilon_{\mu})$. Thus,
$D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}((\epsilon_{\mu}-\epsilon_{\nu})_{+},(\epsilon_{\mu}-\epsilon_{\nu})_{-})\big)^{\frac{1}{p}}\leq D^{1-\frac{1}{p}}\big(W_{1}(\mu,\nu)+W_{1}(\epsilon^{*}_{+},\epsilon^{*}_{-})\big)^{\frac{1}{p}}.$ (29)
A.2 Proof of Theorems
@@
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let ${\epsilon}:G_{n}\to\mathbb{R}$ be a signed noise measure that satisfy Assumption 1. Then, for $p>1$,
$W^{\pm}_{p}(\mu,\mu+{\epsilon})\leq W_{p}({\epsilon}_{-},{\epsilon}_{+}).$ (34)

### v2

@@ 3.3 Multiscale $W_{p}$ bound on a single image and its noisy version
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions. Let $\varepsilon_{1},\varepsilon_{2}$ be independent random signed measures on the grid that satisfy Assumption 1. Then
$\frac{n\sigma}{\sqrt{\pi}}\leq\mathbb{E}W^{\pm}_{1}(\mu+\varepsilon_{1},\mu+\varepsilon_{2})\leq\frac{2\sqrt{2}n\log_{2}n}{\sqrt{\pi}}\sigma+\frac{n}{\sqrt{2\pi}}\sigma.$ (15)
@@
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$. Let $\varepsilon_{1},\varepsilon_{2}$ be independent random signed measures on the grid that satisfy Assumption 1. For convenience, we again assume that $n=2^{\eta}$, for $\eta\in\mathbb{N}$. Then, for $p>1$ with $p\in\mathbb{N}$,
$\mathbb{E}\left[\left(W_{p}^{\pm}(\mu+\varepsilon_{1},\mu+\varepsilon_{2})\right)^{p}\right]\leq\frac{4\sqrt{2}}{\sqrt{\pi}}n\sigma.$ (16)
@@
$\mathbb{E}\left[W_{p}^{\pm}(\mu+\varepsilon_{1},\mu+\varepsilon_{2})\right]\leq\left(\frac{4\sqrt{2}}{\sqrt{\pi}}n\sigma\right)^{1/p}.$
Remark 1.
In both theorems above, the upper bound can be improved by removing the factor $\sqrt{2}$ if only one image is corrupted by noise.
@@ 3.4 Multiscale $W_{p}$ bound on two images and their noisy counterparts
Let $\mu,\nu:G_{n}\to[0,1]$ be two probability measures on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let $\varepsilon_{\mu},\varepsilon_{\nu}:G_{n}\to\mathbb{R}$ be signed noise measures that satisfy Assumption 1. For convenience we assume that $n=2^{\eta}$, for $\eta\in\mathbb{N}$. Then
$\mathbb{E}\left[W^{\pm}_{1}(\mu+\varepsilon_{\mu},\nu+\varepsilon_{\nu})-W^{\pm}_{1}(\mu,\nu)\right]\leq\frac{4n\log_{2}n+n}{\sqrt{\pi}}\sigma+\frac{\sqrt{2}}{n}.$ (18)
The proofs of the three theorems above come from a multiscale upperbound on the Wasserstein distance that depends solely on the mass differences (here, the pixels intensities) and therefore enables to control the impact of the noise.
Even though the Wasserstein 2-distance is often used in applications and has nice theoretical properties in the continuous setting —such as the Brenier–McCann theorem (Brenier1991), its signed counterpart does not enjoy the same metric properties as the signed Wasserstein 1-distance, as was already hinted at in the introduction.
@@
Let $\mu,\nu:G_{n}\to[0,1]$ be two probability measures on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let $\varepsilon_{\mu},\varepsilon_{\nu}:G_{n}\to\mathbb{R}$ be signed noise measures that satisfy Assumption 1. For convenience we assume that $n=2^{\eta}$, for $\eta\in\mathbb{N}$. Then,
$\mathbb{E}\big[W_{p}^{\pm}(\mu+\varepsilon_{\mu},\nu+\varepsilon_{\nu})\big]\leq\left(\frac{\sqrt{2}}{2}\right)^{1-\frac{1}{p}}W_{1}(\mu,\nu)^{\frac{1}{p}}+\frac{\sqrt{2}}{2}\bigg(\frac{4}{\sqrt{\pi}}n\log_{2}n+\frac{2}{\sqrt{\pi}}n\bigg)^{\frac{1}{p}}\sigma^{\frac{1}{p}}.$ (19)
The proof of this theorem requires to first relate the signed Wasserstein $p$ distance to a 1-Wasserstein distance between the uncorrupted measures by exploiting suboptimal couplings. This then enables to use a multiscale bound for the remaining part that mostly pertains to noise.
@@ 4.1 Quantitative validation of noise scaling
The first experiment we conduct aims to quantitatively measure how the distance between an image and its noisy counterpart scales when increasing noise variance. This allows for a direct comparison between the empirical behavior of each metric and the theoretical scaling laws derived in Theorem 3. The results are reported as Figure 2. All transport costs calculated are exact and were calculated by using the POT python package FlamaryCourtyGramfortAlayaBoisbunonEtAl2021
Figure 2: $W_{1},W_{2}$ and $L_{2}$ (markers) distances plotted against their fits and their theoretical bounds from Theorems 2 and 3 (dashed lines).
To this end, we performed 100 independent trials, each time selecting a new, random 32x32 pixel image from the DOTMark 1.0 MicroscopyImages dataset (SchrieberSchuhmacherGottschlich2017). For each image $\mu$, we generated a noisy version $\mu+\varepsilon$ by adding zero-sum noise $\varepsilon$ satisfying Assumption 1, with variances ranging from $10^{-7}$ to $1$. We computed the difference between the original image and the noisy one for $L^{2}$, $W_{2}$ and $W_{1}$ identifying the image with the torus. This empirical result, where the $W_{2}$ distance scales with an exponent of approximately 0.5, suggests that the bound derived in Theorem 3 correctly captures the behavior of the signed 2-Wasserstein as a function of the noise variability $\sigma$.
@@ 4.2 Visualizing robustness of inter-image distances
We now investigate how well the different metrics preserve the original distance between two images when the latter are progressively corrupted by noise. For this experiment, we selected two distinct $32\times 32$ pixel images from the DOTMark dataset and simultaneously corrupted them with different instances of zero-sum additive noise with a standard deviation ranging from $10^{-7}$ to $10^{-1}$. At each noise level, we computed the $W_{1},W_{2}$ and $L^{2}$ distances between the two noisy images. The results were averaged across 100 experiments. To evaluate stability, we computed a “Distance Ratio” by dividing the distance between the noisy images by the constant distance between the original, clean images. A ratio that remains close to 1 indicates that the metric’s measurement reflects the underlying signals rather than the noise. The output is displayed on Figure 1, which we already exhibited in the introduction.
Figure 3: Distances between two randomly sampled images from the DOTMark microscopy dataset, both being noised with noise sampled from the zero-sum normal distribution, in dashed (matching colors) we have the bounds for each $p$ from Theorem 3.
@@
We repeated the distance ratio experiment from the previous section on four distinct image classes: white noise, typical cryo-EM projections, classic microscopy images, and synthetic images of two widely separated squares. These classes represent types of images which are ranging from pure noise, The results are shown in Figure 4. In this figure, $W_{2}$ is shown to scale favorably. However, it also exhibits an peculiar lack of monotonicity w.r.t. the noise level. A phenomenon that we explore in Section 4.4.
4.3 Analysis of cryo electron-microscopy images
Single-particle cryo-electron microscopy (cryo-EM) is a method for reconstructing the 3D structure of proteins and other large molecules. In this method, samples of a molecule of interest are frozen and then imaged using a transmission electron microscopy. This results in many thousands of tomographic projections of the target molecule. The positions and orientations of the individual molecules is typically unknown and the images have extremely high level of noise. Nonetheless, sophisticated computational methods were successful in recovering many different high-resolution 3D structures. Many important challenges remain. In particular, the reconstruction of flexible macromolecules with continuous degrees of freedom. See BendoryBartesaghiSinger2020 for a survey of the computational challenges in cryo-EM. In this section, we wish to demonstrate the potential benefit of Wasserstein metrics in the high-noise cryo-EM regime to the difficult task of recovering continuous conformational manifolds (KileelMoscovichZeleskoSinger2021). To this end, we generated 20 different projections of the E. Coli hsp90 protein (ShiauHarrisSouthworthAgard2006) in different conformational states using cryoJAX (OBrienSilva-SanchezWoollardJeHansonEtAl2025). The location of the protein was shifted and the pictures were corrupted with high levels of noise to mimic the poor signal-to-noise ratio in real cryo electron microscopy images. For simplicity, all the images were normalized to sum to one. The goal is to assess how well each metric can recover known geometric relationships between particle images that undergo rotation and translation. To illustrate the difficulty of the task, Figure 5 shows a sample of the original, clean images alongside their noisy counterparts.
Figure 5: Projection images of the E. Coli hsp90 molecule, with and without noise.
Figure 6: Top panel: Distance matrices between the different structures in the absence of noise. Bottom panel: averaged distances over 100 experiments. A more gradual blue-to-red gradient is better, $W_{2}$ performs best.
@@ 4.4 The decreasing distance phenomenon
Figure 7: Distances between two images, one has a mass of 1 in (8,8) and the other in (24,24) and zero everywhere else, both being noised with noise sampled from the zero-sum normal distribution, in dashed (matching colors) we have the bounds for each $p$ from Theorem 5.
Figure 8: Top - Source image, and where the mass of the original pixel (8,8) goes. Bottom - Target image, and where the mass of the target pixel (24, 24) comes from for the optimal transport map between noised versions of the single pixel images.
@@ 5 Conclusion and future work
Despite these results, there remains venue for additional work. A primary challenge would be to establish a sharp bounds for $\mathbb{E}W_{p}^{\pm}(\mu_{\varepsilon},\nu_{\varepsilon})-\mathbb{E}W_{p}^{\pm}(\mu,\nu)$, which is hindered by the lack of triangle inequality. The numerical experiments further suggest that our bounds, despite capturing the correct behavior, are not tight. Finally, as our theory suggests that robustness increases with higher values of $p$, the interest of such choices of exponents for practical applications should be investigated in future works.
Appendix A Proofs
@@ A.1 Proof of Propositions
Proposition 2
In the context of Assumption 1, the following holds.
  1. 1.
  2. 2.
    The sum of the components is zero : $\sum_{i=1}^{M}N_{i}=0$.
@@
Proposition 3 (Wasserstein Distance Decomposition).
Proof of Proposition 3.
$m(A)\leq\mu(A)\text{ and }m(A)\leq\nu(A).$
The remaining, disjoint parts of each measure are given by $\mu^{\prime}:=\mu-m=(\mu-\nu)_{+}$ as well as $\nu^{\prime}:=\nu-m=(\nu-\mu)_{+}$. Thus, we can write:
$\mu=m+\mu^{\prime}\qquad\nu=m+\nu^{\prime}$ (25)
$W_{p}^{p}(\mu,\nu)\leq W_{p}^{p}(\mu^{\prime},\nu^{\prime})$ (27)
@@
Let $\mu:G_{n}\to[0,1]$ be a probability measure on the $n\times n$ unit grid $G_{n}$ with cyclic boundary conditions and let $\varepsilon:G_{n}\to\mathbb{R}$ be a signed noise measure that satisfy Assumption 1. Then, for $p>1$,
$W^{\pm}_{p}(\mu,\mu+\varepsilon)\leq W_{p}(\varepsilon_{-},\varepsilon_{+}).$ (28)